OpenStudy (anonymous):

Four friends are going to sit together in a roller-coaster car that has 4 seats. In how many different arrangements can they sit?

OpenStudy (anonymous):

24 ways

OpenStudy (anonymous):

It's 4!.. ( 4 Factorial permutaion problem)

OpenStudy (anonymous):

\[\huge \color{green}{= ^{4}P_4 = \frac{4!}{(4-4)!} = \frac{4!}{1!} = 4!}\]